### v2

# TULIP: Transformer for Upsampling of LiDAR Point Clouds

Abstract
LiDAR Upsampling is a challenging task for the perception systems of robots and autonomous vehicles, due to the sparse and irregular structure of large-scale scene contexts. Recent works propose to solve this problem by converting LiDAR data from 3D Euclidean space into an image super-resolution problem in 2D image space. Although their methods can generate high-resolution range images with fine-grained details, the resulting 3D point clouds often blur out details and predict invalid points. In this paper, we propose TULIP, a new method to reconstruct high-resolution LiDAR point clouds from low-resolution LiDAR input. We also follow a range image-based approach but specifically modify the patch and window geometries of a Swin-Transformer-based network to better fit the characteristics of range images. We conducted several experiments on three different public real-world and simulated datasets. TULIP outperforms state-of-the-art methods in all relevant metrics and generates robust and more realistic point clouds than prior works. We plan to publish the code upon acceptance.
$\dagger$\$$\dagger$\$footnotetext: Authors share last authorship.
1 Introduction
(a) Input
(b) Ground Truth
(c) SwinIR [26]
(d) TULIP (Ours)
Figure 1: TULIP generates more realistic LiDAR point clouds from low-resolution LiDAR input (a) compared to the existing image super-resolution approach [26].
Light Detection And Ranging (LiDAR) is one of the most common sensors for perception in various fields of autonomy, such as autonomous driving, and unmanned aerial vehicles (UAV). LiDARs are used to generate 3D point clouds of the scene. These point clouds are essential for mapping, localization, and object detection tasks. However, the accuracy of these tasks often depends on the resolution of the point cloud [48]. Furthermore, the resolution of a LiDAR is inherently associated with increased energy consumption and cost, making its use impractical for various applications. LiDARs also have different vertical and horizontal resolutions. The vertical resolution of rotating 3D LiDARs is typically much lower than the horizontal resolution. These limitations make an upsampling technique necessary to increase the resolution of the LiDAR data, especially in the vertical direction. Moreover, upsampling LiDAR data has the potential to counter domain shift problems in LiDAR-based learning methods. Notably, the use of lower resolution LiDAR data in a system that has been trained on higher resolution data leads to a significant drop in performance [3]. Therefore, such upsampling techniques can not only help to mitigate the domain gap by creating a virtual high-resolution sensor that matches the target domain [37, 56] but also reduce the high cost of collecting new LiDAR data, annotating and retraining the methods.
Deep learning has led to remarkable advancements in LiDAR upsampling techniques in recent years. Several recent methods [51, 53, 37] have focused on learning the upsampling process within the 3D Euclidean space. However, the processing of 3D data in deep learning can be quite resource-intensive. One approach to this is the representation of the point cloud as a range image [38, 24, 17, 44]. This approach facilitates the adoption of well-established image-based super-resolution methods. Image-based super-resolution methods mostly employ autoencoder-style networks with convolutional and deconvolutional layers [38, 57, 13]. However, these methods are not directly transferable to range images. Range images generated by back projection of sparse 3D point clouds have distinct sharp edges at the object boundaries. Convolutional neural network based methods often induce edge smoothing due to the regularization effect, making them unsuitable for LiDAR range image processing.
As an alternative, Transformer [43] has shown great success in RGB image super-resolution and producing less blurry outputs [26]. Vision transformers, however, face challenges such as the need for large amounts of training data [9, 26, 31] and the global computation of self-attention, limiting the capture of local information in the range images. To address this, Swin Transformer [29] introduced tokenizing the input image and applying the self-attention mechanism locally in independent windows. While the modifications to Swin Transformer have led to robust training and promising results compared to Vision Transformer, its default structure and settings are not ideally suited for processing LiDAR range images. These range images consist of single-channel data, representing 3D spatial information, as opposed to three-channel color images that depict visual appearance. Furthermore, range images feature large, smooth areas within 3D objects, separated by distinct sharp edges between objects. As a result, minor inaccuracies in the 2D prediction (e.g. blurry output) can lead to significant differences in the projected 3D occupancy. This non-uniform distribution of image details further complicates network training. The challenging details in the range image are concentrated in a few specific pixels, contrasting with RGB images where relevant details are evenly distributed. This makes range images fundamentally different from RGB images. Inspired by the limitations of state-of-the-art range image upsampling methods and the potential of Swin-Transformer [29], we propose a novel network named TULIP. Our proposed geometry-aware architecture is tailored to accommodate LiDAR range image data. The proposed method takes a low-resolution range image as input and produces a high-resolution range image, which can then be projected into a 3D point cloud. To accommodate LiDAR range image geometry, we utilize one-line row patches to tokenize the input range image, as opposed to the square patches used for RGB images. This approach helps to preserve vertical information for the upsampling process while capturing boundary discontinuities effectively. This also creates a trade-off between local detail capture and model complexity in the training process. Additionally, we incorporate a non-square window for local self-attention computation. This enhances learning of spatial contexts at different scales in the range image. Further, we extensively train and evaluate TULIP on two real-world and one simulated autonomous driving datasets: KITTI [19], DurLAR [25] and CARLA [24], respectively. Comparisons to existing approaches demonstrate that TULIP outperforms the state-of-the-art methods on all three datasets.
2 Related Works
Image super-resolution has seen significant progress in recent years. The technique aims to construct high-resolution (HR) images from low-resolution (LR) observations, often leveraging advances in Convolutional Neural Networks (CNNs) to enhance the fidelity and detail of images for better visualization and information extraction. Subsequently, several enhanced frameworks have been developed [13, 14, 54, 23, 22].
Driven by their success in the field of natural language processing (NLP) [21, 4], Transformers  [43] has been extended to solve a variety of vision-related tasks such as object detection [6] and semantic segmentation [2, 7]. Such Vision Transformers (ViT) [16] excels at learning to focus on relevant image regions by exploring global interactions among different regions. Their impressive performance has led to their adoption in image restoration tasks [26, 9, 11] as well. While they brought significant improvements in RGB image super-resolution, Vision Transformer [16] comes with the drawback of the quadratic computational complexity of self-attention. Additionally, they capture mostly global dependencies within the data and require large amounts of data for training.
To address these challenges, Swin Transformer [29] has been proposed. Unlike the Vision Transformer, which relies on global self-attention across the entire image, the Swin Transformer employs a local window-based attention mechanism and establishes a pyramid-like architecture. This approach processes images hierarchically by gradually merging smaller image patches into larger ones. This strategy enables more effective handling of different scales and facilitates the processing of multi-scale features, making it particularly suitable for range image processing. Furthermore, due to Swin Transformer’s hierarchical structure, the increase in the number of model parameters scales linearly with the input image size, which mitigates the challenges associated with handling large range images. Additionally, they require less training data than the classical Vision Transformer, which is beneficial for LiDAR upsampling as LiDAR datasets are typically magnitudes smaller than RGB datasets used for Vision Transformer. The Swin and Vision Transformer have also been used for image super-resolution on omnidirectional panoramic camera images [40, 50]. However, these approaches are specifically designed for the unique characteristics of these cameras (e.g. strong distortions, spherical continuity) that are vastly different from the LiDAR sensing model, which renders them unsuitable for application to range images.
Other related works have focussed on upsampling high-density point clouds from low-density point clouds  [56, 51, 49, 55, 34]. While those approaches also provide qualitative results for LiDAR point clouds, they focus on increasing the general 3D point density of point clouds of single objects instead of a large scene. Another area closely associated with LiDAR upsampling is depth completion [12, 20, 47, 41, 35, 33, 32, 27]. The primary objective of depth completion is to improve sparse depth estimates acquired from LiDAR by integrating data from multiple sensors, predominantly RGB cameras. The result of depth completion methods is a dense depth map that provides depth values to every pixel in the input depth map. Nonetheless, this field slightly diverges from LiDAR upsampling due to its reliance on multimodal sensors and its common deployment within a restricted field of view, making it not directly comparable within the scope of this paper. LiDAR point clouds exhibit specific characteristics, such as a distinct point pattern (due to the stacked lasers in rotating 3D LiDARs) and a decrease in point density with increasing distance from the sensor. Differently from arbitrary point cloud upsampling, LiDAR upsampling tries to mimic a realistic point cloud of a high-resolution LiDAR given the point cloud of low-resolution LiDAR and thus targets a different result than the approaches above.
Due to the difficulty and high computational cost of upsampling LiDAR point clouds in 3D space [37, 1, 8], most works  [24, 38, 42] represent the point cloud as a range image and perform the LiDAR upsampling task in 2D image space. LiDAR-CNN [42] introduces a CNN-based architecture with semantic and perceptual guidance. The method specifies two further loss functions besides a per-point reconstruction loss to acquire a better synthesis of the high-resolution range image. However, this approach is limited to data with semantic annotations, which drastically reduces its applicability. LiDAR-SR [38] deploys a CNN-based network with U-Net [36] architecture. It additionally uses a Monte-Carlo dropout post-processing to reduce the amount of noisy points in the prediction. These approaches based on convolutional operations tend to fail to reconstruct the sharpness in the range image effectively. While the previous approaches directly predict the range image, Implicit LiDAR Network (ILN) [24] uses an implicit neural architecture that learns interpolation weight to fill in new pixels instead of their values directly. Although the method outperforms CNN-based approaches in training speed and preserving the geometrical details in the input, it can still suffer from limited neighboring information in the LiDAR data, especially in areas distant from the sensor due to the extremely sparse input data. Differently from the previous works, our approach builds on Swin-Transformer as a backbone and is specifically customized to effectively process range images.
3 Methodology
LiDAR upsampling is achieved by upsampling a range image, effectively transforming it from a 3D upsampling problem into a 2D image super-resolution problem [24, 42, 38, 17]. It is, therefore, an evident approach to build on RGB image super-resolution. However, image super-resolution aims to solve a different problem than LiDAR upsampling. Image super-resolution enhances the visual appearance of a low-resolution RGB image by recreating a high-resolution image. While quantitative metrics for evaluating these works exist, there is no one specific correct solution. Differently, LiDAR upsampling tries to recreate a high-resolution, 3D LiDAR sensor output from a low-resolution LiDAR input point cloud. Range images contain more geometrical and spatial contexts rather than visual information, and the scene’s geometry strictly dictates the correct solution. The underlying sensing model is also drastically different, as LiDARs are active sensors. The respective range pixel values purely depend on distance, while pixels in an RGB camera depend on many factors, such as scene appearance, lighting conditions, exposure time, and white balance. Furthermore, range images have a highly asymmetric field-of-view (FoV), between 1:8 to 1:64, which differs drastically from regular camera images. Accordingly, most works [42, 17, 38] only perform vertical upsampling, in contrast to image super-resolution, which operates in both dimensions. A naive application of state-of-the-art image super-resolution networks [26] to range images does not result in adequate performance (Fig. 1). To this end, we developed TULIP, a novel range image-based LiDAR Upsampling method. Based on the observations above, we specifically adapt a network based on SwinUnet [5] to incorporate range images better. In the following sections, we will describe the technical details of our method, with a special focus on how our approach accommodates range images.
3.1 Problem Definition
Figure 2: Using a spherical projection, the point cloud (top) can be represented as the range image (bottom) and vice versa.
A LiDAR point cloud consists of points captured during one revolution $\mathcal{P}=\{\mathbf{p}_{1},,...,\mathbf{p}_{n}\}$. Each measurement represent at 3D point $\mathbf{p}_{i}=\{x_{i},y_{i},z_{i}\}$. As input, we assume a low-resolution point cloud $\mathcal{P}_{l}$ with $n_{l}=H_{l}\times W_{l}$ points, where $H_{l}$ and $W_{l}$ correspond to vertical and horizontal resolution. We aim to predict a high-resolution point cloud $\mathcal{P}_{h}$, that has the same FoV as $\mathcal{P}_{l}$ but contains $n_{h}=\beta*n_{l}$ points. In our experiments, we set $\beta=4$. We only increase the point cloud’s vertical resolution (amount of LiDAR active beams), i.e., $n_{h}=H_{h}\times W_{h}$, where $H_{h}=\beta*H_{l}$ and $W_{h}=W_{l}$. We project the point cloud into a 2D range image. In this image, each row and column coordinate $v,u$ correspond to the respective elevation and azimuth angles of the LiDAR points, while the pixel value contains the range of the point $r=\sqrt{x^{2}+y^{2}+z^{2}}$. The image coordinate of a 3D point can be calculated using a spherical projection model:
$\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}\frac{W}{2}-\frac{W}{2\pi}arctan(\frac{y}{x})\\ \frac{H}{\Theta_{max}-\Theta_{min}}*(\Theta_{max}-arctan(\frac{z}{\sqrt{x^{2}+y^{2}}})\end{bmatrix}$ (1)
$\Theta_{max}$ and $\Theta_{min}$ correspond to the limits of the vertical field of view of the LiDAR. We formulate the LiDAR upsampling problem as an image upsampling problem, i.e. we aim to predict a high-resolution range image $I_{h}\in\mathbf{R}^{1\times H_{h}\times W}$ given a low-resolution range image $I_{l}\in\mathbf{R}^{1\times H_{l}\times W}$. We can then calculate the high-resolution point cloud $\mathcal{P}_{h}$ by inverting the projection in Eq. 1.
3.2 Architecture Design and Overview
Our network builds upon Swin-Unet [5] which is originally designed for image segmentation. As shown in Fig. 3, we deploy a U-shaped network structure featuring skip connections that link the encoder and decoder modules. The fundamental building block of TULIP is the Swin Transformer [29]. TULIP first uses a convolutional layer to patchify and map the input to a high-dimensional feature space. The tokenized low-level feature maps are passed through a multi-stage encoder. Each stage contains a pair of Swin Transformer [29] blocks and a patch merging layer to downsample the resolution by a factor of two and increase the dimension by a factor of four. This step realizes the hierarchic computation and extraction of multi-scale features through encoding. Within each stage, multiple instances of the self-attention mechanism are performed locally in parallel, which is so-called Window Multi-Head Self Attention (W-MSA). After passing through a two-layer MLP and residual connection with the input, the feature vector is further passed to the second part of the Swin Transformer block, utilizing a Shifted Window Multi-Head Self Attention (SW-MSA) which extends W-MSA by employing a shifted window operation. This enhances the model capacity as it compensates for the lack of interaction between the local windows in W-MSA. The bottleneck layer, which consists of two subsequent Swin Transformer [29] blocks, is applied to enhance the high-level feature representation further. The decoder, which realizes the upsampling has a symmetric design with respect to the encoder and operates in a reverse way of the encoder. The resolution of the feature maps is first expanded, and the dimension is reduced by a factor of two accordingly. We feed existing geometrical information via skip connections. The feature maps are transformed into a single-channel, high-resolution range image in the last layer. The projection head comprises a $1\times 1$ convolutional layer for feature expansion, followed by a Leaky ReLU activation, a pixel shuffle layer [39], and another $1\times 1$ convolutional layer for final projection. As the loss function, we select the pixel-wise L1 loss.
Figure 3: Network architecture of TULIP: We adapt the network presented in [5] for range image upsampling. The low-resolution input is encoded using Swin Transformer blocks and patch merging layers (left). Patch unmerging layers are used to rearrange the feature patches for upsampling (left). The projection head generates the high-resolution range image from the latent features.
3.2.1 Patch Partition
The patch partition layer creates an initial feature embedding from the input range image. This initial feature representation significantly influences the network’s performance. Specifically, the selection of this layer determines how the inherent characteristics, patterns, and relevant information of the LiDAR range image data are encoded into a format that the network can effectively learn from. Most transformer-based RGB image super-resolution approaches utilize a relatively large patch size to tokenize the input [26, 9, 31]. These larger patches enable the construction of a global spatial context, facilitating an understanding of the overall structure and high-level features of the RGB image. These networks employ square-shaped patches and implement upsampling in both vertical and horizontal directions. In contrast, LiDAR upsampling is primarily aiming to enhance the vertical resolution of the input, given that range images possess properties that are vastly different from RGB images. Furthermore, attention across more pixels in range image is less useful than in RGB image, since there is almost no geometrical relation for the spatial contexts that are far from each other in the 2D range image space. Motivated by this, we propose two adjustments to effectively process range images within a Swin-Unet:
Row-Based Patch Tokenization: Our model builds row patches with a dimension of $1\times 4$ for range image tokenization. The new patch geometry is designed to retain full vertical information while compressing horizontally. This aligns with our objective of conserving and extending vertical details in the range image. Besides that, row patches excel at capturing boundary discontinuities between distinct objects and the background within the scene. Moreover, the selection of the patch size strikes a balance between enhancing model performance and limiting model capacity. Smaller patches can capture more detailed local information, but they also increase the number of model parameters, which can slow down the model’s training process. Circular Padding (CD) in Horizontal Dimension: Instead of adding zero pixels padding, our model uses circular padding in the horizontal dimension. This avoids introducing artificial features and naturally matches the sensor model of a rotating 3D LiDAR. This preserves accurate neighbor relations along the edges, which would otherwise be compromised during the projection to the 2D range image. The Circular Padding (CD) also allows us to fit flexible input sizes without changing the patch geometry.
3.2.2 Non-Square window for local self-attention
Inside Swin Transformer, input feature maps are window-partitioned along width and height to compute the self-attention locally in each window. In addition to square-shaped RGB images, some recent work applies a square window for panoramic data as well [28, 52]. However, using square windows one range images is unfavorable. On one hand, small windows can focus more on attention within near areas but lead to increasing computational complexity and can struggle with the reconstruction of objects at a further distance in the scene. On the other hand, a larger square window can blend scene-related contexts at different scales, which can potentially degrade the network performance. In a prior study done by Eskandar et al. [17], they demonstrated the significant enhancement in range image super-resolution by separately upsampling the upper and lower portions of the data. In contrast to their approach of utilizing CNN-based shallower and deeper network branches for distinct feature extraction, our proposal involves the direct partitioning of the input using a rectangular window. This window geometry facilitates robust attention interaction among distant points through W-MSA and subtle cross-attention between objects at multiple scales via SW-MSA and hierarchical processing (refer to Fig. 4). In addition, range images, as typical panorama data, inherently contain abundant information along the horizontal direction. Therefore, directing attention more toward the image width rather than the height can aid the network in capturing a greater amount of geometrical information. At the same time, the computational complexity remains the same as using a large square window.
Figure 4: Our network uses a hierarchical structure for patch merging on the range image. We propose to use rectangle window to fit the range data and to handle depth discontinuity at different scales.
3.2.3 Additional Modifications and Postprocessing
Patch Unmerging (PU) and Pixel Shuffle (PS): Similar to the patch merging layer that is responsible for downsampling at the encoder stage, a mechanism is necessary within the decoder to upsample the feature maps. Uformer [45] achieves upsampling through a de-convolutional layer. As they tend to smooth out image sharpness [26], we design a reverse operation of patch merging. We re-arrange the channels of patches into their respective positions in a grid to upsample the resolution of input feature maps, which we denote as patch unmerging. Furthermore, we built the projection head based on the pixel shuffle operation [39]. Monte Carlo Dropout: We additionally use Monte Carlo Dropout (MC-Dropout) [38] to filter unreliable points. MC-Dropout executes several feed-forward passes during inference with different active dropouts, which yields a distribution of outputs. We refine the results by removing points with high output variance, as they often come down to noisy, invalid points.
4 Experiments
4.1 Experimental Settings
Datasets: We conduct experiments on three different datasets, that include the two large-scale real-world datasets DurLAR [25] and KITTI [19] (Sec. 4.3.2), as well as on a dataset [24] that was generated using the CARLA simulator [15] (Sec. 4.3.1). For all datasets, we select test sequences that are recorded in different locations than the train set to avoid spatial overlap. We vertically downsample the high-resolution range images with a factor of four by skipping the respective lines to generate low-resolution input images. For ablation studies in Sec. 4.2, we select the KITTI dataset. Implementation Details: In all experiments, we use distributed processing with 4×4$\times$4 × GeForce RTX 2080 Ti. For training, we use a fixed batch size of eight per GPU for all datasets. We use the AdamW [30] optimizer with a weight decay factor of 0.05 and a base learning rate of 5e-4. Evaluation metrics: We evaluate the Mean Absolute Error (MAE) for all the pixels in the generated 2-D range images. Additionally, we assess the performance by considering the 3D points reconstructed by our neural networks. The Chamfer Distance (CD) measures the Euclidean distance between two point clouds. We also follow the approach in [24] to evaluate the volumetric occupancy similarity. To do so, the point clouds are voxelized using a voxel size of $0.1m$. A voxel is classified as occupied for each point cloud if it contains at least one point. We then calculate the Intersection-over-Unit (IoU) based on the occupancy.
4.2 Ablation Studies
Model Patch Window CP/PU/PS MC MAE $\downarrow$ IoU $\uparrow$ CD $\downarrow$
Baseline [5] $4\times 4$ Square $\times$ $\times$ 0.7138 0.3250 0.1940
Model 1 $2\times 2$ Square $\times$ $\times$ 0.6251 0.3337 0.1661
Model 2 $1\times 4$ Square $\times$ $\times$ 0.4814 0.3667 0.1391
Model 3 $1\times 4$ Rectangular $\times$ $\times$ 0.4228 0.4040 0.1218
Model 4 $1\times 4$ Rectangular $\checkmark$ $\times$ 0.4186 0.4094 0.1199
TULIP $1\times 4$ Rectangular $\checkmark$ $\checkmark$ 0.4183 0.4134 0.1198
TULIP-L $1\times 4$ Rectangular $\checkmark$ $\checkmark$ 0.3678 0.4329 0.0971
Table 1: Ablation study results for 4×4$\times$4 × upsampling on KITTI. Input Resolution: $16\times 1024$, Output Resolution: $64\times 1024$.
Patch size: In Tab. 1 Baseline and Model 1-2, we compare the upsampling results with different patch sizes. We observe that the smaller patch ($2\times 2$) outperforms the larger one ($4\times 4$) while the proposed row patch ($1\times 4$) can further refine the network performance significantly. In particular, the row patch improves $32.6\%$ in MAE, $12.8\%$ in IoU, and $28.3\%$ in CD, compared to a $4\times 4$ patch. As visualized in Fig. 5, the row patch better preserves sharpness around edges and corners compared to the larger patch sizes.
(a) Scene
(b) GT
(c) $1\times 4$
(d) $2\times 2$
(e) $4\times 4$
Figure 5: Upsampling results with different patch sizes: We observe that square patches ($4\times 4$, $2\times 2$) blur out edges and create invalid connections between separate walls, while the proposed $1\times 4$ patch generates sharp edges and correct discontinuities.
(a) Scene
(b) GT
(c) Rectangular
(d) Square
Figure 6: Upsampling results with different window geometries: Rectangular windows improve reconstruction of objects at different ranges, indicated by more distinct discontinuities (top) and a better reconstruction of the car (bottom).
Window geometry: To verify the effectiveness of modifying window geometry for range images, we conduct experiments with square and rectangular windows (Tab. 1 Model 2-3. Based on the quantitative results, we can observe the following: reconstruction quality produced by local attention within large square windows is inferior to the rectangular window with respect to both 2D and 3D evaluation metrics, while they result in the same computational complexity. The underlying reason for this performance difference lies in the perceptive field of the windows. A wider window is more likely to include discontinuities between objects at different scales. Accordingly, the network performs better at separating objects from each other, which can visually be validated by clearer surface boundaries in Fig. 6.
Additional Modifications and Postprocessing: We additionally apply Circular Padding (CP), Patch Unmerging (PU), and Pixel Shuffle (PS) layer in Model 4. Due to space restrictions and a smaller impact on the performance, we evaluate them together. For TULIP, we also apply Monte Carlo dropout [18] as a post-processing step, which helps to further reduce ghost points in between objects, which is reflected by the enhancement in IoU. For TULIP-L, we increase the model capacity, using four Swin-Transformer layers instead of three in both encoder and decoder, which demonstrates further performance boost on all metrics.
Inference Time: We evaluate the inference time on the KITTI dataset [19]. A single forward pass for TULIP-L takes 1.33s. For the Monte-Carlo Dropout we run the separate forward passes in a batch and calculate the filtered result, which increases the effective inference time of TULIP-L to 1.61s. This is slower than ILN [24] (1.14s) and LIDAR-SR [38] (1.16s) and comparable to Swin-IR [26] (1.63s).
4.3 Benchmark Results
Besides bilinear interpolation, which computes the interpolation weights from the four nearest neighbors for upsampling, we compete against the state-of-the-art LiDAR upsampling approaches Implicit LiDAR Network (ILN) [24] and LiDAR-SR [38]. We also compare against several image super-resolution works: SRNO  [46], and LIIF [10] learn implicit features for interpolation instead of 2D coordinates of new pixels while HAT [9] and Swin-IR [26] are pixel-based approaches that directly reconstruct high-resolution images from low-resolution feature embeddings. Similar to our work, they are also both based on Swin Transformer [29]. We train all methods on the specific datasets using their default parameters.
CARLA 4x Output Resolution: $128\times 2048$
Model MAE $\downarrow$ IoU $\uparrow$ CD $\downarrow$
Bilinear 1.8128 0.1382 0.7262
SRNO [46] 2.4640 0.1343 2.0230
HAT [9] 1.6032 0.2698 0.6337
SWIN-IR [26] 1.9560 0.2718 0.4840
LIIF [10] 0.8064 0.3502 0.174
LIDAR-SR [38] 0.8216 0.2581 0.2044
ILN [24] 0.8592 0.5206 0.1855
TULIP (Ours) 0.7768 0.5291 0.1168
TULIP-L (Ours) 0.7748 0.5332 0.1032
Table 2: Quantitative comparison against state-of-the-art LiDAR and image super-resolution methods on simulation dataset. All methods are trained and evaluated on the same splits.
KITTI 4x Output Resolution: $64\times 1024$
Model MAE $\downarrow$ IoU $\uparrow$ CD $\downarrow$
Bilinear 2.0892 0.1063 0.6000
SRNO [46] 0.8350 0.2035 0.4417
HAT [9] 0.6856 0.2035 0.2516
SWIN-IR [26] 1.2972 0.2774 0.7347
LIIF [10] 0.6143 0.3226 0.1916
LIDAR-SR [38] 0.5674 0.1005 0.2165
ILN [24] 1.0528 0.3342 0.2787
TULIP (Ours) 0.4183 0.4134 0.1198
TULIP-L (Ours) 0.3678 0.4329 0.0971
DurLAR 4x Output Resolution: $128\times 2048$
Bilinear 2.4384 0.1266 0.6346
SRNO [46] 1.5396 0.1507 0.5108
HAT [9] 1.7820 0.2353 0.1973
SWIN-IR [26] 1.9416 0.2157 0.2279
LIIF [10] 1.5672 0.2469 0.1548
LIDAR-SR [38] 1.5312 0.1370 0.1128
ILN [24] 1.5720 0.3430 0.0893
TULIP (Ours) 1.5382 0.3552 0.06560
TULIP-L (Ours) 1.5386 0.3620 0.06371
Table 3: Quantitative comparison against state-of-the-art LiDAR and image super-resolution methods on real-world dataset. All methods are trained and evaluated on the same splits.
Input
Swin-IR [26]
LIIF [10]
LiDA-SR [38]
ILN [24]
TULIP (Ours)
GT
A
B
C
Figure 7: Qualitative results on KITTI: We provide visualizations for the top-4 results from Tab. 3 and Swin-IR [26]. Our approach outperforms other state-of-the-art methods in upsampling realistic point clouds. A: TULIP is the only approach that does not generate invalid ghost points in between the car (right bottom) and the wall behind. It can also be observed that all approaches except for ILN match the characteristic line pattern of the LiDAR on the ground (middle). B: Swin-IR and LIIF generate a blurry reconstruction. TULIP achieves the clearest point cloud and even reconstructs small details such as the car’s side mirrors. C: Swin-IR, LIIF, and LiDAR-SR do not clearly separate the car from the wall behind. ILN generates a clear point cloud but does not properly reconstruct the shape of the cars. TULIP achieves the clearest point cloud, with distinct discontinuities between objects, and closely resembles the shape of the three cars.
4.3.1 Simulation Results
Similar to prior works, we first evaluate the performance on noise-free, simulated data. We train and test all approaches on the dataset presented in  [24] that was recorded using the CARLA simulator and select the ground-truth range images of size $128\times 2048$ that capture a vertical FoV of ${}^{\circ}$. We use a (20699/2618) train/test split. We observe that the performance between image super-resolution approaches varies greatly, and LIIF performs the best among them. While LIIF [10] achieves a comparable MAE to the LiDAR upsampling methods, the 3D metrics are significantly worse, indicating that it is unsuitable for the underlying geometric input data. Nevertheless, LIIF [10] still outperforms the LiDAR upsampling method in LIDAR-SR. The interpolation-based approach in ILN shows decent results for the noise-free data, as indicated by their high IoU. Nevertheless, our approaches achieve the best performance in all evaluated metrics, with the strongest improvement in Chamfer Distance, which validates that our design is well suited for the geometric reconstruction in range images.
4.3.2 Real-World Results
To evaluate our approach on real-world data, we downsample the DurLAR and KITTI datasets to a comparable number of frames to the CARLA dataset by temporally skipping frames in the dataset sequences. In particular, we have the following (train/test) split: DurLAR (24372/2508), and KITTI (20000/2500).
KITTI: The KITTI dataset [19] was collected using a Velodyne HDL-64E LiDAR with a vertical FoV of ${}^{\circ}$ and a resolution of $64\times 1024$. For the image super-resolution approaches, we observe comparable results and trends on KITTI as on the simulated data in Sec.4.3.1. Even though LIDAR-SR [38] is designed for LIDAR upsampling and outperforms image super-resolution approaches in terms of MAE, it achieves the lowest IoU among all approaches, which can be explained by a large amount of invalid floating points between real objects. LIIF [10] achieves better CD and comparable IoU compared to ILN. We observe that the IoU of ILN [24] is drastically lower than on CARLA. The interpolation scheme in ILN is thus not suitable to properly handle the lower resolution data with noise from the real-world sensor. Our approaches significantly outperform all competing methods, especially in terms of IoU as it can better handle discontinuities between objects. As indicated in the visualizations in Fig.7, our approach predicts fewer invalid float points between objects, as well as fewer blurry edges, which results in much clearer point clouds.
DurLAR: The DurLAR dataset [25] was recorded using an Ouster OS1-128 LiDAR, that captures point clouds with a resolution of $128\times 2048$ with a ${}^{\circ}$ vertical FoV. Compared to CARLA and KITTI, we see a large performance drop for all approaches in MAE and IoU, which results from the larger maximum range and vertical FoV of the sensor as well as stronger sensor noise. We observe that all approaches designed for LiDAR upsampling clearly improve in Chamfer Distance compared to image super-resolution approaches. ILN and TULIP outperform all methods by a large margin in IoU. TULIP-L achieves the best 3D metrics, which validates that it can also handle the challenging DurLAR data better than prior works.
Range Analysis: As the density of LiDAR point clouds decreases with increasing distance from the sensor, we additionally compare the performance at different ranges for a more fine-grained evaluation. We observe that our approach achieves better or similar performance compared to all approaches in all metrics at all ranges. Up to 30m, TULIP clearly outperforms the baseline approaches. TULIP handles discontinuities between objects better than other approaches, which can be verified visually through sharper edges and fewer ”floating” points between objects. This is especially reflected in the high IoU. At ranges above 30m, our approach performs comparable to other methods. Above this range, LiDAR data is extremely sparse, meaning that points neighboring in the low-resolution range image are actually distant in the Euclidean space. It is, therefore, not feasible to infer meaningful information about the scene from such neighboring points at a high range, which is also represented in the poor performance of all approaches.
(a) IoU $\uparrow$
(b) Chamfer distance $\downarrow$
Figure 8: 3D error metrics visualized at different ranges for KITTI.
4.4 Failure Cases
Although TULIP outperforms other state-of-the-art methods both qualitatively and quantitatively, the upsampling quality can still be limited in certain cases. For instance, as shown in Fig 9, our approach shows inferior upsampling results for the specific scene. The irregularity of the scene leads to high uncertainty, which makes the reconstruction noisy. Observing the details in Fig. 8(a), the network fails to reconstruct the car at the top of the scene.
(a) Full scene
(b) Details
Figure 9: Failure Case: noisy point cloud generated by TULIP and missing reconstruction of the object. (Ours and GT)
5 Conclusion
This work presents TULIP, a novel method for LiDAR Upsampling that achieves incredible performance in upsampling the range image. Our approach transforms 3D point clouds into 2D range images and performs the upsampling in 2D space. We build upon a Swin-Transformer-based network and specifically modify the patch partition and attention windows to better accommodate the characteristics of range images. Throughout various experiments, testing on three different benchmarks, it shows that TULIP outperforms state-of-the-art methods quantitatively up to 35.3% in MAE, 49.3% in Chamfer distance, and 29.5% in IoU. Quantitatively, TULIP demonstrates more promising results in upsampling the realistic LiDAR data.